### 01 Dec 2023 v2

@@ 6 Application in polyphase code design
We approach (25) for $n=4,5,\ldots,11$ with the 5th order relaxations of the three hierarchies. Related results are shown in Table 7. We make the following observations: (1) for $n=4,5,6$, R-HSOS, C-HSOS, and R-SOS provide the same lower bound; (2) for $n\geq 7$, R-HSOS and C-HSOS provide the same lower bound while R-SOS runs out of memory; (3) R-HSOS is the fastest.
@@
$n$ opt R-HSOS C-HSOS R-SOS
4 0.5000 0.01 0.02 2.04
5 1.0000 0.02 0.04 22.6
6 4.0000 0.06 0.18 497
7 1.1418 0.21 0.38 -
8 1.7428 0.57 1.94 -
9 0.0594 2.74 8.95 -
10 3.4932 13.4 68.7 -
11 2.2096 80.4 401 -
@@
We approach (27) for $n=4,5,\ldots,15$ with the 4th order relaxations of the three hierarchies. Related results are shown in Table 8. We make the following observations: (1) for $n=4,5,6,7$, R-HSOS, C-HSOS, and R-SOS provide the same lower bound; (2) for $n\geq 8$, R-HSOS and C-HSOS provide the same lower bound while R-SOS runs out of memory; (3) R-HSOS is the fastest.
@@
$n$ opt R-HSOS C-HSOS R-SOS
4 0.5000 0.01 0.02 1.43
5 0.7703 0.02 0.04 10.5
6 1.0000 0.05 0.12 73.2
7 0.5219 0.14 0.33 732
8 0.6483 0.37 0.77 -
9 0.1119 1.03 2.67 -
10 0.5805 2.30 7.21 -
11 0.4943 6.37 24.2 -
12 0.4928 19.3 87.8 -
13 0.4189 105 440 -
14 0.3309 324 1548 -
15 0.3098 1096 5546 -

### 25 Feb 2024 v3

@@ 6 Application in polyphase code design
We approach (25) for $n=4,5,\ldots,12$ with the 5th order relaxations of the three hierarchies. The related results are shown in Table 7 where the upper bounds computed by Ipopt are also displayed for comparison. We make the following observations: (1) for $n=4,5,6$, R-HSOS, C-HSOS, and R-SOS provide the same lower bound matching the corresponding upper bound; (2) for $n\geq 7$, R-HSOS and C-HSOS provide the same lower bound matching the corresponding upper bound up to $n=10$ while R-SOS runs out of memory; (3) R-HSOS is the fastest.
@@
$n$ ub R-HSOS C-HSOS R-SOS
opt time opt time opt time
4 0.5000 0.5000 0.01 0.5000 0.02 0.5000 2.04
5 1.0000 1.0000 0.02 1.0000 0.04 1.0000 22.6
6 4.0000 4.0000 0.06 4.0000 0.18 4.0000 497
7 1.1418 1.1418 0.21 1.1418 0.38 - -
8 1.7428 1.7428 0.57 1.7428 1.94 - -
9 0.0594 0.0594 2.74 0.0594 8.95 - -
10 3.4932 3.4932 13.4 3.4932 68.7 - -
11 2.2579 2.2096 80.4 2.2096 401 - -
12 5.7178 -0.7769 350 - - - -
@@
We approach (27) for $n=4,5,\ldots,15$ with the 4th order relaxations of the three hierarchies. The related results are shown in Table 8 where the upper bounds computed by Ipopt are also displayed for comparison. We make the following observations: (1) for $n=4,5,6,7$, R-HSOS, C-HSOS, and R-SOS provide the same lower bound matching the corresponding upper bound; (2) for $n\geq 8$, R-HSOS and C-HSOS provide the same lower bound matching the corresponding upper bound up to $n=9$ while R-SOS runs out of memory; (3) R-HSOS is the fastest.
@@
$n$ ub R-HSOS C-HSOS R-SOS
opt time opt time opt time
4 0.5000 0.5000 0.01 0.5000 0.02 0.5000 1.43
5 0.7703 0.7703 0.02 0.7703 0.04 0.7703 10.5
6 1.0000 1.0000 0.05 1.0000 0.12 1.0000 73.2
7 0.5219 0.5219 0.14 0.5219 0.33 0.5219 732
8 0.6483 0.6483 0.37 0.6483 0.77 - -
9 0.1119 0.1119 1.03 0.1119 2.67 - -
10 0.8248 0.5805 2.30 0.5805 7.21 - -
11 0.6671 0.4943 6.37 0.4943 24.2 - -
12 0.8971 0.4928 19.3 0.4928 87.8 - -
13 0.7154 0.4189 105 0.4189 440 - -
14 0.8146 0.3309 324 0.3309 1548 - -
15 0.9232 0.3098 1096 0.3098 5546 - -